# Online Decision-Focused Learning

### v2

@@ Results.
We believe that this work can be extended in several ways. First, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators. Second, it would be valuable to investigate less adversarial environments, such as those involving i.i.d. data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge. Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical perspective.
Appendix A H˜2 is satisfied in the Gaussian linear case.
@@
As a concrete example, consider $\mathbf{O}$ as the stochastic gradient descent (SGD) algorithm. Even in the context of deep networks, it is plausible that $\mathbf{O}$ converges to a local minimizer. Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [31] or almost surely [55, 58, 20]. Such stationary points may correspond to local/global minima or saddle points.
Recent studies further show that SGD avoids saddle points. For instance, Mertikopoulos et al. [55] proved that the trajectories of SGD almost surely avoid all strict saddle manifolds—i.e., sets of critical points $x$ where the Hessian has at least one negative eigenvalue. These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [48].
Beyond SGD, similar guarantees extend to more general methods. In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [40].
Appendix C Supplementary results on the simplex.
@@ C.1 Additional framework
This setting encompasses various decision-making scenarios such as portfolio selection [47] or mixed strategy design in multiplayer games [68]. In this case, we choose the negative entropy
@@ Appendix D Background on the FTPL algorithm
The recent work of [67] proposed an online algorithm for nonconvex losses $(\ell_{t})_{t\in[T]}$ with static regret guarantees. They rely on an approximate optimization oracle $\mathcal{O}$ which takes as input any function $\ell$, a $d$-dimensional vector $\sigma$ and returns an approximate minimizer of $\ell-(\sigma,\cdot)$. An optimization oracle is called "( $\xi,\chi$ )-approximate optimization oracle" if it returns $\mu^{*}\in\mathcal{K}$ such that
@@
Given access to an ($\xi,\chi$ )-approximate optimization oracle, [67] study the FTPL algorithm which is described by the following recursion. Starting from $\hat{\mu}_{1}$, at each time steps $t$, $\hat{\mu}_{t}\in\mathcal{K}$ is updated as follows:
@@
Proposition 4 (Theorem 1 of [67]).
@@
$\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
@@ Appendix E Additional experimental material
The motivating example from [54].
@@
Figure 2: Figure 2 from [54].
@@
Appendix F Regret bound with an approximate optimal solution.
@@ Appendix G Proofs
G.1 Proofs of preliminary lemmas.
@@ G.1.1 Simplex
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)\;.$
@@
We now proceed to bound $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ for $\beta\in(0,1)$. With $a=\alpha_{t}\ln(d)$ and $b=2\ln(d)$, we have:
$\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}]\;,$
and since for any $x>0,\,1-x\leqslant e^{-x}\leqslant 1-x+x^{2}$:
$\varphi(\beta)$ $\leqslant(1-\beta)^{-1}[1-(1-\beta)\ln(a)+(1-\beta)^{2}\ln^{2}(a)-1+(1-\beta)\ln(b)]$
$=\ln\mathopen{}\mathclose{{\left(\frac{b}{a}}}\right)+(1-\beta)\ln^{2}(a)\;.$
Substituting in the values of $a$ and $b$ yields the desired result. ∎
@@ G.1.2 General polytope
we have $h_{t}(w)\rightarrow+\infty$ as $w\rightarrow\text{bdry}(\mathcal{W}_{1})$. Since $h_{t}(\tilde{w}_{t}(\theta),\theta)=\min_{w\in\mathcal{W}_{1}}h_{t}(w,\theta)$, we deduce that $\tilde{w}_{t}(\theta)\in\text{int}(\mathcal{W}_{1})$. It follows from the first order condition and the implicit function theorem [21, Theorem 2] that $\tilde{w}_{t}:\theta\mapsto\tilde{w}_{t}(\theta)$ is differentiable, and
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
@@
which is uniquely determined by strong convexity of the objective. We know by [14, page 566] that:
@@ G.1.3 Other lemmas
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (39)
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$ (40)
To bound the first term in the right-hand-side of Equation˜40. by the definition of conditional expectation, we have:
@@ G.3 Proof of Theorem˜2
$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)\;.$
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{\left(1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)+\xi T\;.$ (46)
@@
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;.$ (48)
@@ G.4 Proof of Proposition˜2
Then, taking the $\sup$ over $\Theta$ in Equations˜49 and 51, and defining
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (52)
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (53)

### v3

@@ 5 Conclusion
We believe that this work can be extended in several ways. First, it is plausible that faster rates (e.g., $T^{-1/2}$) could be obtained using more traditional online-learning techniques that bypass the bilevel framework, such as discretizing the parameter space and applying expert aggregation over the resulting bins. However, such approaches would likely incur a prohibitive dependence on the parameter dimension $m$ (e.g., exponential in $m$). Exploring this direction remains however an interesting avenue in future work. Second, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operators, to adress the more challenging of a general convex set $\mathcal{W}$. Third, it would be valuable to investigate less adversarial environments, such as those involving i.i.d. data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge. Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical perspective.
Acknowledgment
Funded by the European Union (ERC, Ocean, 101071601). M.H. is partially supported by the European Research Council Starting Grant DYNASTY – 101039676. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A H˜2 is satisfied in the Gaussian linear case
@@
As a concrete example, consider $\mathbf{O}$ as the stochastic gradient descent (SGD) algorithm. Even in the context of deep networks, it is plausible that $\mathbf{O}$ converges to a local minimizer. Indeed, a large body of work has analyzed SGD in non-convex settings, establishing convergence to stationary points either in expectation [Ghadimi and Lan, 2013] or almost surely [Mertikopoulos et al., 2020, Patel and Zhang, 2021, Cutkosky et al., 2023]. Such stationary points may correspond to local/global minima or saddle points.
Recent studies further show that SGD avoids saddle points. For instance, Mertikopoulos et al. [2020] proved that the trajectories of SGD almost surely avoid all strict saddle manifolds—i.e., sets of critical points $x$ where the Hessian has at least one negative eigenvalue. These manifolds include connected families of non-isolated saddle points, a phenomenon common in the loss landscapes of overparametrized neural networks [Li et al., 2018].
Beyond SGD, similar guarantees extend to more general methods. In particular, the stochastic Riemannian Robbins–Monro method (a broad template encompassing various algorithms) has been shown to converge almost surely to local or global minima [Hsieh et al., 2023].
Finally, we highlight that, in most practical industrial settings, solving the downstream optimization problem is typically far more computationally demanding than updating the prediction model. It is therefore reasonable to assume that the oracle call represents only a small fraction of the overall computational cost.
Appendix C Supplementary results on the simplex
@@ C.1 Additional framework
This setting encompasses various decision-making scenarios such as portfolio selection Li and Hoi [2014] or mixed strategy design in multiplayer games Syrgkanis et al. [2015]. In this case, we choose the negative entropy
@@ Appendix D Background on the FTPL algorithm
The recent work of Suggala and Netrapalli [2020] proposed an online algorithm for nonconvex losses $(\ell_{t})_{t\in[T]}$ with static regret guarantees. They rely on an approximate optimization oracle $\mathcal{O}$ which takes as input any function $\ell$, a $d$-dimensional vector $\sigma$ and returns an approximate minimizer of $\ell-(\sigma,\cdot)$. An optimization oracle is called "( $\xi,\chi$ )-approximate optimization oracle" if it returns $\mu^{*}\in\mathcal{K}$ such that
@@
Given access to an ($\xi,\chi$ )-approximate optimization oracle, Suggala and Netrapalli [2020] study the FTPL algorithm which is described by the following recursion. Starting from $\hat{\mu}_{1}$, at each time steps $t$, $\hat{\mu}_{t}\in\mathcal{K}$ is updated as follows:
@@
Proposition 4 (Theorem 1 of Suggala and Netrapalli [2020]).
@@
$\frac{1}{T}\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)}}\right)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
@@ Appendix E Additional experimental material
The motivating example from Mandi et al. [2024].
@@
Figure 2: Figure 2 from Mandi et al. [2024].
@@
Higher dimension problem.
Finally, the following figure shows the performance of our algorithms in a higher dimension problem, with $K=80$ items. We see that our algorithms still significantly beat the baselines, suggesting that they remain effective in moderately high dimension problems.
Figure 4: Experiment results with 80 items.
Appendix F Regret bound with an approximate optimal solution
@@ Appendix G Proofs
G.1 Proofs of preliminary lemmas
@@ G.1.1 Simplex
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}+2(\alpha_{t}\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$
@@
We now proceed to bound $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ for $\beta\in(0,1)$. With $a=\alpha_{t}\ln(d)$ and $b=2\ln(d)$, we have
$\varphi(\beta)=(1-\beta)^{-1}(a^{\beta-1}-b^{\beta-1})=(1-\beta)^{-1}\bigl[e^{-(1-\beta)\ln(a)}-e^{-(1-\beta)\ln(b)}\bigr]\;.$
Factoring out $e^{-(1-\beta)\ln(a)}=a^{\beta-1}$ yields
$\varphi(\beta)=a^{\beta-1}\,(1-\beta)^{-1}\Bigl[1-e^{-(1-\beta)\ln(b/a)}\Bigr]\;.$
Since for any $x\in\mathbb{R}$ one has $1-e^{-x}\leqslant\min\{x,1\}$, we obtain
$\varphi(\beta)\leqslant a^{\beta-1}\,(1-\beta)^{-1}\min\Bigl\{(1-\beta)\ln(b/a),\,1\Bigr\}=a^{\beta-1}\min\Bigl\{\ln(b/a),\,(1-\beta)^{-1}\Bigr\}\;.$
Substituting the values of $a$ and $b$ finally gives
$\varphi(\beta)\leqslant(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;,$
which concludes the proof.
@@ G.1.2 General polytope
we have $h_{t}(w)\rightarrow+\infty$ as $w\rightarrow\text{bdry}(\mathcal{W}_{1})$. Since $h_{t}(\tilde{w}_{t}(\theta),\theta)=\min_{w\in\mathcal{W}_{1}}h_{t}(w,\theta)$, we deduce that $\tilde{w}_{t}(\theta)\in\text{int}(\mathcal{W}_{1})$. It follows from the first order condition and the implicit function theorem [de Oliveira, 2014, Theorem 2] that $\tilde{w}_{t}:\theta\mapsto\tilde{w}_{t}(\theta)$ is differentiable, and
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}$
@@
which is uniquely determined by strong convexity of the objective. We know by [Boyd, 2004, page 566] that:
@@ G.2 Proof of Theorem˜1
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$ (39)
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)\;.$ (40)
We now bound the first term in the right-hand-side of Equation˜40. By the definition of conditional expectation, we have:
@@
$+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha(\alpha\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)$
@@ G.3 Proof of Theorem˜2
Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (42) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (44) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant|\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]|\leqslant\sup_{\theta\in\Theta}|\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]|\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[|f_{t}(\theta)-\tilde{f}_{t}(\theta)|]$ leads to:
$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right|}}\right]+\xi T$
@@
$\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]\leqslant\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}\;.$
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\mathopen{}\mathclose{{\left(\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)+\xi T\;.$ (46)
@@
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathopen{}\mathclose{{\left(\alpha_{t}+2n\max_{k\in[K]}\mathopen{}\mathclose{{\left\|v_{k}}}\right\|_{1}\alpha_{t}(\alpha_{t}\ln d)^{\beta-1}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)+\xi\;.$ (48)
@@ G.4 Proof of Proposition˜2
Then, taking absolute values, using the triangular inequality and taking the $\sup$ over $\Theta$ in Equations˜49 and 51, as well as defining
@@
$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right|}}\right]$
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha+2(\alpha\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}\;,\;.$ (52)
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}T\mathopen{}\mathclose{{\left(\alpha+2(\alpha\ln(d))^{\beta}C_{0}\min\Bigl\{\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right),\,(1-\beta)^{-1}\Bigr\}}}\right)\;.$ (53)